# UL-UNAS: Ultra-Lightweight U-Nets for Real-Time Speech Enhancement via Network Architecture Search

### v1

@@ Abstract
Lightweight models are essential for real-time speech enhancement applications. In recent years, there has been a growing trend toward developing increasingly compact models for speech enhancement. In this paper, we propose an Ultra-Lightweight U-Net optimized by Network Architecture Search (UL-UNAS), which is suitable for implementation in low-footprint devices. Firstly, we explore the application of various efficient convolutional blocks within the U-Net framework to identify the most promising candidates. Secondly, we introduce two boosting components to enhance the capacity of these convolutional blocks: a novel activation function named affine PReLU and a causal time-frequency attention module. Furthermore, we leverage neural architecture search to discover an optimal architecture within our carefully designed search space. By integrating the above strategies, UL-UNAS not only significantly outperforms the latest ultra-lightweight models with the same or lower computational complexity, but also delivers competitive performance compared to recent baseline models that require substantially higher computational resources.
keywords:
speech enhancement , ultra-lightweight , neural architecture search , computational complexity
journal: Neural Networks
@@
[…]
1 Introduction
Speech enhancement (SE) aims to recover clean speech from the noisy mixture, thereby improving speech quality and intelligibility. In recent years, significant breakthroughs in SE have been achieved due to the rapid evolution of deep neural networks (DNNs). In general, DNN-based SE algorithms can be categorized into the time-frequency (T-F) domain approaches (Tan and Wang, 2018; Hu et al., 2020; Le et al., 2021; Yu and Luo, 2023; Maier, 2022; Wang et al., 2023) and the time-domain approaches (Pandey and Wang, 2019; Luo and Mesgarani, 2019; Luo et al., 2020; Wang et al., 2021). For the T-F domain approaches, the short-time Fourier transform (STFT) representation of noisy speech is fed into a DNN, which outputs an enhanced speech spectrogram through techniques such as masking (Tan and Wang, 2018; Hu et al., 2020; Le et al., 2021; Yu and Luo, 2023), deep filtering (Maier, 2022; Mack and Habets, 2019), or mapping (Wang et al., 2023). Phase estimation is a tricky problem in these cases as the phase spectrum usually presents an irregular structure due to the wrapping effect. Earlier T-F domain-based works only focus on the recovery of spectral magnitude by using the ideal ratio mask (IRM) (Narayanan and Wang, 2013), leaving the phase unaltered. In contrast, phase-sensitive mask (Erdogan et al., 2015) takes into account the cosine of the phase difference between clean and noisy speech when estimating IRM and shows improved results. Complex ideal ratio mask (cIRM) (Williamson et al., 2015) applies a complex instead of a real gain to perform implicit phase correction, leading to better perceptual quality but negligible improvements in objective intelligibility. Directly predicting the real and imaginary parts of the clean spectrum can also implicitly recover the phase information, and usually has a higher upper bound of performance than that of masking methods (Tan and Wang, 2019; Wang et al., 2023).
Compared to the T-F domain approaches, time domain approaches circumvent the challenging phase estimation problem by conducting feature encoding and decoding in a trainable manner. They can achieve very low algorithm latency due to the learnable signal analysis and synthesis on very short windows. However, they sacrifice the advantages of STFT, which provides a feasible and robust feature representation where speech and noise components exhibit distinct structural characteristics in the spectral domain. Consequently, the performance of time domain methods is not among the top tier in SE tasks (Yin et al., 2020). Differently, recent multi-domain methods, which combine the advantages of both the T-F domain and time domain approaches to build a composite SE system, often exhibit better performance than single-domain methods (Fan et al., 2023; Wang and Wang, 2022).
Although DNN-based approaches have achieved overwhelming performances over traditional SE algorithms, their improved performances are often accompanied by large computation overhead. Most state-of-the-art (SOTA) SE models require substantial computational resources, making them infeasible for deployment on edge devices for practical applications. Therefore, exploring lightweight SE neural networks with performance comparable to the SOTA models is increasingly becoming a research hotspot. One straightforward solution is model compression, including pruning (Stamenovic et al., 2021), quantization (Tan and Wang, 2021), and inference skipping (Le et al., 2022). TinyLSTMs (Fedorov et al., 2020) applies these three techniques to a simple long short-term memory (LSTM) SE network, achieving a 2.9$\times$ reduction in operations with minor performance degradation.
@@ 1 Introduction
Another commonly adopted approach is to hand-craft efficient architectures. TRU-Net (Choi et al., 2021) replaces the standard convolution with depthwise separable convolution and employs a more powerful dual-path recurrent neural network (RNN) structure. In Braun et al. (2021), it is demonstrated that parallel RNN grouping and convolutional additive skip connections are highly effective in the design of compact architectures, enabling the corresponding model to stand out within the family of manually crafted architectures. Furthermore, combining the advantages of traditional signal processing and DNNs has demonstrated remarkable potential. RNNoise (Valin, 2018) and PercepNet (Valin et al., 2020) first perform a coarse enhancement on a low-resolution spectral envelope and then apply finer noise suppression between pitch harmonics using a pitch comb filter. Similarly, DeepFilterNet (Schröter et al., 2022; Maier, 2022) performs enhancement on compact ERB features while reconstructing periodic components through deep filtering utilizing coefficients predicted from both ERB and the original STFT features. DPCRN-CF (Le et al., 2023) utilizes a DNN-based pitch estimator and a learnable comb filter to achieve superior harmonic enhancement. Despite the significant reduction in computational overhead achieved by these works, they are still too large for practical deployment on edge devices with low power consumption requirements, except for RNNoise, which is compact enough but suffers from limited performance.
@@
On the other hand, neural architecture search (NAS) is an approach to automating the model design process. NAS aims to find the optimal architecture based on a specific performance estimation strategy within the defined search space, using an appropriate search strategy (Elsken et al., 2019). NASNet (Zoph et al., 2018) employs a controller network to sample architecture candidates from a certain search space. Subsequently, each candidate is trained and evaluated to obtain accuracy as a reward. Eventually, the controller is updated through a reinforcement learning (RL) algorithm. MnasNet (Tan et al., 2019) goes further by introducing a novel factorized search space, along with taking latency into account when computing reward, thereby facilitating the design of a platform-aware low-latency network. EfficientNet (Tan and Le, 2019) employs the same NAS strategy as MnasNet with the exception that it focuses on optimizing floating-point operations (FLOPs) instead of latency, thereby preventing itself from being restricted to any particular hardware device.
@@
The number of parameters and computational complexity, represented by floating-point operations per second (FLOPS) or multiply-accumulate operations per second (MACS), serve as the principal guidelines for researchers in the design of lightweight models. The number of parameters determines the memory footprint of models and holds particular significance when the objective is to deploy models on memory-constrained devices. However, in most cases, it is latency, rather than memory, that acts as the bottleneck of edge devices. In such cases, computational complexity can offer a more pertinent and useful metric than the number of parameters. When comparing different architectures, the number of parameters does not provide highly accurate or detailed insights, as different architectures typically exhibit varying levels of parameter efficiency (Zhang et al., 2024). For instance, a fully connected network can have massive parameters yet still manage to achieve a relatively fast inference speed. In contrast, a transformer network may have fewer parameters but run much slower due to the computationally expensive attention operations. For models of the same architecture type, increasing the number of parameters can lead to improved performance. ParameterNet (Han et al., 2024) demonstrates this in the context of large-scale visual pretraining. By leveraging dynamic convolution to add parameters with only a marginal increase in FLOPS, it achieves superior results.
@@
In this paper, we focus on optimizing the ultra-lightweight SE model with computational complexity as low as around 30M MACS while imposing no strict constraints on the number of parameters. We adopt the T-F domain approach due to its performance advantage under low computational resource constraints. Building upon our previously proposed ultra-lightweight SE model, GTCRN (Rong et al., 2024), we conduct a comprehensive exploration of the encoder-decoder design to improve both model capacity and efficiency. Furthermore, by incorporating NAS, we refine the U-Net architecture, ultimately achieving an optimal configuration. Our contributions are summarized as follows:
  • 1.
@@
  • 4.
    Our proposed UL-UNAS, which has a smaller computational complexity, outperforms existing SOTA ultra-lightweight SE models. With only 34M MACS and a compact size of 169k parameters, UL-UNAS achieves a PESQ score of 3.09 on the VCTK-DEMAND dataset. This advancement paves the way for more efficient and effective ultra-lightweight design in various application scenarios.
2 Related Work
@@
2.1 Ultra-lightweight speech enhancement models
@@
Over the past few years, the concept of ultra-lightweight SE neural networks has gradually emerged, aiming to be compatible with edge devices with extremely limited computing resources. GTCRN (Rong et al., 2024) represents one of the seminal efforts in creating ultra-lightweight SE models. It attains competitive performance compared to several typical larger models while requiring no more than 40M MACS. Another ultra-lightweight SE model named FSPEN (Yang et al., 2024) was proposed during the same period, featuring a computational complexity of 89M MACS. LiSenNet (Yan et al., 2024), which requires a computational cost of 56M MACS, achieves the highest performance among them.
@@
All the aforementioned models are built upon the U-Net backbone thanks to its computationally efficient encoder-decoder (ED) architecture, while incorporating modifications to improve both performance and efficiency. The modifications are centered around two key components: the ED structure and the bottleneck enhancer. Regarding the ED component, GTCRN incorporates grouped convolution along with lightweight boosting modules including sub-band feature extraction and lightweight attention to strengthen the basic building blocks; FSPEN utilizes a dual-ED architecture comprising sub-band and full-band processing for better speech feature extraction; LiSenNet introduces novel sub-band downsampling and upsampling blocks specifically designed for band-aware feature capture. As for the bottleneck enhancer, GTCRN adopts a parallel RNN grouping strategy to reduce computational burden; FSPEN develops a path extension method to improve inter-frame modeling; LiSenNet introduces a gate mechanism at the end of the dual-path architecture to augment model capacity. Our work inherits the advantageous designs in the GTCRN architecture and focuses on exploring more efficient designs for the building blocks within the ED structure.
2.2 Efficient designs of convolutional blocks
@@
Numerous innovative concepts have been introduced to design efficient convolutional networks to improve their efficacy. According to Ma et al. (2024), these insights can be divided into several categories. Among them, depthwise separable convolution (Howard, 2017; Sandler et al., 2018; Howard et al., 2019) decomposes a standard convolution into a depthwise and a pointwise convolution. Feature shuffling (Zhang et al., 2018; Ma et al., 2018) simplifies pointwise convolution by group convolution and channel shuffle. Feature re-use (Han et al., 2020) generates additional feature maps through inexpensive linear transformation, enabling more efficient information exploitation. Reparameterization (Vasu et al., 2023) introduces multiple parallel branches in training to improve performance, which are merged into one branch before inference. Star operation (Ma et al., 2024) augments the representative ability of a convolutional block by mapping inputs into a high-dimensional non-linear feature space. In this paper, we explore the potential applications of these designs in ultra-lightweight SE neural networks, as detailed in Sec. 3.2.1.
2.3 Nonlinear activation functions
@@
The design of nonlinear activation functions has been an active area of research in neural networks, as it directly impacts the capabilities of deep learning models. The early proposed smooth nonlinear activation functions such as Sigmoid and Tanh, suffer from the problem of vanishing gradients, which constrains the performance of DNNs. The rectified linear unit (ReLU) (Nair and Hinton, 2010), by presenting a non-saturating nonlinearity, alleviates this issue and has emerged as a preferred option in modern networks. To tackle the “dying ReLU” problem, some variants have been proposed, such as leaky ReLU (Maas et al., 2013) and parametric ReLU (PReLU) (He et al., 2015), by permitting a small, non-zero gradient when the input is negative. Utilizing a weighted combination of base activation functions has also been explored. Adaptive piecewise linear activation unit (Agostinelli et al., 2014) computes the activation output through a weighted sum of multiple ReLU units with learnable weights and bias. VanillaNet (Chen et al., 2023) introduces a series-informed activation function to enhance the nonlinearity of a neural network. It achieves this by concurrently stacking the activation functions, which can be regarded as a series in a mathematical sense. To further enrich the approximation ability of the series, the spatial neighbors of a specific element of the inputs are taken into account to produce outputs. However, this causes it to function in a manner analogous to a depthwise convolution operation, leading to an increase in computational complexity. In contrast, our proposed APReLU (as described in Sec. 3.2.2) integrates a learnable affine transformation with PReLU, creating a more powerful activation function with negligible computational overhead.
2.4 Efficient attention modules
@@
Attention modules are found effective in boosting the performance of neural networks. By pooling the feature maps along non-relevant dimensions to obtain a reduced representation that retains global information, the model is able to attend to the salient features of the target dimension. This process usually requires a relatively low computation cost due to the dimension reduction. SENet (Hu et al., 2018) presented a novel channel-attention block that adaptively recalibrates channel-wise feature responses based on input features pooled over spatial dimensions. CBAM (Woo et al., 2018) goes further by incorporating a cascaded spatial attention module based on input features pooled over channel dimensions. TFA (Zhang et al., 2022) introduces a similar attention mechanism in SE, achieving a T-F distribution modeling by two parallel attention branches operating on features pooled over the frequency and time dimensions, respectively. However, the average pooling along the time axis within the frequency attention branch results in non-causality, thus hindering its application in real-time SE models. Although the TRA introduced in GTCRN overcomes the non-causality issue by discarding the frequency attention branch, it sacrifices the potential benefits of frequency-dependent attention. In contrast, our proposed cTFA (as described in Sec. 3.2.3) preserves this frequency-dependent branch while effectively addressing the non-causality problem by diverting around the time-axis pooling.
2.5 Neural architecture search for speech enhancement

### v2

@@ Abstract
Lightweight models are essential for real-time speech enhancement applications. In recent years, there has been a growing trend toward developing increasingly compact models for speech enhancement. In this paper, we propose an Ultra-Lightweight U-Net optimized by Network Architecture Search (UL-UNAS), which is suitable for implementation in low-footprint devices. Firstly, we explore the application of various efficient convolutional blocks within the U-Net framework to identify the most promising candidates. Secondly, we introduce two boosting components to enhance the capacity of these convolutional blocks: a novel activation function named affine PReLU and a causal time-frequency attention module. Furthermore, we leverage neural architecture search to discover an optimal architecture within our carefully designed search space. By integrating the above strategies, UL-UNAS not only significantly outperforms the latest ultra-lightweight models with the same or lower computational complexity, but also delivers competitive performance compared to recent baseline models that require substantially higher computational resources. Source code and audio demos are available at https://github.com/Xiaobin-Rong/ul-unas.
I Introduction
Speech enhancement (SE) aims to recover clean speech from the noisy mixture, thereby improving speech quality and intelligibility. In recent years, significant breakthroughs in SE have been achieved due to the rapid evolution of deep neural networks (DNNs). In general, DNN-based SE algorithms can be categorized into the time-frequency (T-F) domain approaches [47, 18, 21, 64, 32, 58] and the time-domain approaches [35, 27, 26, 57]. For the T-F domain approaches, the short-time Fourier transform (STFT) representation of noisy speech is fed into a DNN, which outputs an enhanced speech spectrogram through techniques such as masking [47, 18, 21, 64], deep filtering [32, 31], or mapping [58]. Phase estimation is a tricky problem in these cases, as the phase spectrum usually presents an irregular structure due to the wrapping effect. Earlier T-F domain-based works only focused on the recovery of spectral magnitude by using the ideal ratio mask (IRM) [34], leaving the phase unaltered. In contrast, phase-sensitive mask [6] takes into account the cosine of the phase difference between clean and noisy speech when estimating IRM and shows improved results. Complex ideal ratio mask (cIRM) [59] applies a complex instead of a real gain to perform implicit phase correction, leading to better perceptual quality but negligible improvements in objective intelligibility. Directly predicting the real and imaginary parts of the clean spectrum can also implicitly recover the phase information, and usually has a higher upper bound of performance than that of masking methods [48, 58].
Compared to the T-F domain approaches, time domain approaches circumvent the challenging phase estimation problem by conducting feature encoding and decoding in a trainable manner. They can achieve very low algorithm latency due to the learnable signal analysis and synthesis on very short windows. However, they sacrifice the advantages of STFT, which provides a feasible and robust feature representation where speech and noise components exhibit distinct structural characteristics in the spectral domain. Consequently, the performance of time domain methods is not among the top tier in SE tasks [63].
@@ I Introduction
Another commonly adopted approach is to hand-craft efficient architectures. TRU-Net [4] replaces the standard convolution with depthwise separable convolution and employs a more powerful dual-path recurrent neural network (RNN) structure. In [2], it is demonstrated that parallel RNN grouping and convolutional additive skip connections are highly effective in the design of compact architectures, enabling the corresponding model to stand out within the family of manually crafted architectures. Furthermore, combining the advantages of traditional signal processing and DNNs has demonstrated remarkable potential. RNNoise [55] and PercepNet [54] first perform a coarse enhancement on a low-resolution spectral envelope and then apply finer noise suppression between pitch harmonics using a pitch comb filter. Similarly, DeepFilterNet [44, 32] performs enhancement on compact ERB features while reconstructing periodic components through deep filtering utilizing coefficients predicted from both ERB and the original STFT features. DPCRN-CF [23] utilizes a DNN-based pitch estimator and a learnable comb filter to achieve superior harmonic enhancement. Despite the significant reduction in computational overhead achieved by these works, they are still too large for practical deployment on edge devices with low power consumption requirements, except for RNNoise, which is compact enough but suffers from limited performance.
@@
On the other hand, neural architecture search (NAS) is an approach to automating the model design process. NAS aims to find the optimal architecture based on a specific performance estimation strategy within the defined search space, using an appropriate search strategy [5]. NASNet [68] employs a controller network to sample architecture candidates from a certain search space. Subsequently, each candidate is trained and evaluated to obtain accuracy as a reward. Eventually, the controller is updated through a reinforcement learning (RL) algorithm. MnasNet [50] goes further by introducing a novel factorized search space, along with taking latency into account when computing reward, thereby facilitating the design of a platform-aware low-latency network. EfficientNet [51] employs the same NAS strategy as MnasNet with the exception that it focuses on optimizing floating-point operations (FLOPs) instead of latency, thereby preventing itself from being restricted to any particular hardware device.
@@
The number of parameters and computational complexity, represented by floating-point operations per second (FLOPS) or multiply-accumulate operations per second (MACS), serve as the principal guidelines for researchers in the design of lightweight models. The number of parameters determines the memory footprint of models and holds particular significance when the objective is to deploy models on memory-constrained devices. However, in most cases, it is latency, rather than memory, that acts as the bottleneck of edge devices. In such cases, computational complexity can offer a more pertinent and useful metric than the number of parameters. When comparing different architectures, the number of parameters does not offer particularly informative insights, as different architectures typically exhibit varying levels of parameter efficiency [66]. For instance, a fully connected network can have massive parameters yet still manage to achieve a relatively fast inference speed. In contrast, a transformer network may have fewer parameters but run much slower due to the computationally expensive attention operations. For models of the same architecture type, increasing the number of parameters can lead to improved performance. ParameterNet [9] demonstrates this in the context of large-scale visual pretraining. By leveraging dynamic convolution to add parameters with only a marginal increase in FLOPS, it achieves superior results.
@@
In this paper, we focus on optimizing the ultra-lightweight SE model with computational complexity as low as around 30M MACS while imposing no strict constraints on the number of parameters. We adopt the T-F domain approach due to its performance advantage under low computational resource constraints. Building on our previously proposed ultra-lightweight SE model, GTCRN [41], we further explore convolutional encoder-decoder (CED) designs to improve both capacity and efficiency of the U-Net architecture. In addition, we apply NAS to refine its structure and identify an optimal configuration. Our contributions are summarized as follows:
@@
  • Our proposed UL-UNAS, with lower computational complexity, outperforms existing SOTA ultra-lightweight SE models. With only 35M MACS and a compact size of 171k parameters, UL-UNAS achieves a PESQ score of 3.09 on the VCTK-DEMAND dataset. This advancement paves the way for more efficient and effective ultra-lightweight designs in various application scenarios.
II Related Work
@@
II-A Ultra-lightweight speech enhancement models
@@
Over the past few years, the concept of ultra-lightweight SE neural networks has gradually emerged, aiming to be compatible with edge devices with extremely limited computing resources. Several seminal ultra-lightweight SE models include GTCRN [41] (34M MACS), FSPEN [62] (89M MACS), and LiSenNet [61] (56M MACS). Among these, LiSenNet achieves the highest performance while maintaining a relatively low computational cost.
@@
All the aforementioned models are built upon the U-Net backbone with a dual-path bottleneck enhancer, thanks to its computational efficiency. Various modifications have been introduced into this base architecture to further enhance performance. The modifications are centered around two key components: the CED structure and the bottleneck enhancer. Regarding the CED component, GTCRN incorporates grouped convolution along with lightweight boosting modules, including sub-band feature extraction and lightweight attention, to strengthen the basic building blocks. FSPEN utilizes a dual-CED architecture comprising sub-band and full-band processing for better speech feature extraction. LiSenNet introduces novel sub-band downsampling and upsampling blocks specifically designed for band-aware feature capture. As for the bottleneck enhancer, GTCRN adopts a parallel RNN grouping strategy to reduce computational burden. FSPEN develops a path extension method to improve inter-frame modeling. LiSenNet introduces a gate mechanism at the end of the dual-path architecture to augment model capacity. Given that GTCRN exhibits the lowest computational cost and best matches our target (around 30M MACS), we select it as the backbone and focus on exploring more efficient designs for the building blocks within the CED structure.
II-B Efficient designs of convolutional blocks
@@
Numerous innovative concepts have been introduced to design efficient convolutional networks to improve their efficacy. According to [29], these insights can be divided into several categories. Among them, depthwise separable convolution [15, 43, 16] decomposes a standard convolution into a depthwise and a pointwise convolution. Feature shuffling [67, 28] simplifies pointwise convolution by group convolution and channel shuffle. Feature re-use [10] generates additional feature maps through inexpensive linear transformation, enabling more efficient information exploitation. Reparameterization [56] introduces multiple parallel branches in training to improve performance, which are merged into one branch before inference. Star operation [29] augments the representative ability of a convolutional block by mapping inputs into a high-dimensional non-linear feature space. In this paper, we explore the potential applications of these designs in ultra-lightweight SE neural networks, as detailed in Sec. III-C1.
II-C Nonlinear activation functions
@@
The design of nonlinear activation functions has been an active area of research in neural networks, as it directly impacts the capabilities of deep learning models. The early proposed smooth nonlinear activation functions, such as Sigmoid and Tanh, suffer from the problem of vanishing gradients, which constrains the performance of DNNs. The rectified linear unit (ReLU) [33], by presenting a non-saturating nonlinearity, alleviates this issue and has emerged as a preferred option in modern networks. To tackle the “dying ReLU” problem, in which standard ReLU units may become inactive when their inputs fall entirely within the negative region, some variants have been proposed, such as leaky ReLU [30] and parametric ReLU (PReLU) [13], by permitting a small, non-zero gradient when the input is negative. Utilizing a weighted combination of base activation functions has also been explored. Adaptive piecewise linear activation unit [1] computes the activation output through a weighted sum of multiple ReLU units with learnable weights and bias. VanillaNet [3] introduces a series-informed activation function to enhance the nonlinearity of a neural network. It achieves this by concurrently stacking the activation functions, which can be regarded as a series in a mathematical sense. To further enrich the approximation ability of the series, the spatial neighbors of a specific element of the inputs are taken into account to produce outputs. However, this causes it to function in a manner analogous to a depthwise convolution operation, leading to an increase in computational complexity. In contrast, our proposed APReLU (as described in Sec. III-C2) integrates a learnable affine transformation with PReLU, creating a more powerful activation function with negligible computational overhead.
II-D Efficient attention modules
@@
Attention modules are found to be effective in boosting the performance of neural networks. They typically operate by aggregating information across selected dimensions to produce a compact, context-rich representation, enabling the model to focus on salient features along the target dimension. This process typically incurs low computational cost due to dimensionality reduction. SENet [17] aims to boost the representational capacity of networks by explicitly modeling interdependencies between channels, introducing a novel channel-attention block that adaptively recalibrates channel-wise feature responses based on input features pooled over spatial dimensions. CBAM [60] extends this idea by also capturing spatial interdependencies, incorporating a cascaded spatial attention module based on input features pooled over channel dimensions. TFA [65] introduces a similar attention mechanism for speech enhancement by modeling time and frequency interdependencies in parallel, achieving a T-F distribution modeling by two parallel attention branches operating on features pooled over the frequency and time dimensions, respectively. However, the average pooling along the time axis within the frequency attention branch results in non-causality, thus hindering its application in real-time SE models. Although the TRA introduced in GTCRN overcomes the non-causality issue by discarding the frequency attention branch, it sacrifices the potential benefits of frequency-dependent attention. In contrast, our proposed cTFA (as described in Sec. III-C3) preserves this frequency-dependent branch while offering an alternative, lightweight mechanism for modeling frequency dependencies without relying on time-axis pooling.
II-E Neural architecture search for speech enhancement